### v1

@@ E.3 Approximation
Appendix F Scaling Law
F.1 Generalization Bound of Empirical Risk Minimizer
Remark F.1.
@@ F.1 Generalization Bound of Empirical Risk Minimizer
Lemma F.3 (Lemma 11 of [63], modified by Theorem 2.1 in [36] and in the setting of this paper).
@@
$\|F-F^{*}\|_{L^{F}({\cal X})}^{2}\leq 4\inf_{F^{\prime}\in{\cal F}_{\sf M,N}(\mathbb{D})}\|F-F^{*}\|_{L^{F}({\cal X})}^{2}+L{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}d}\cdot O(\frac{(B_{F}^{2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+\xi^{2})}V_{({\cal F}_{\sf M,N}(\mathbb{D}),{\rm d})}(\varepsilon)}{n}+(B_{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+\xi)}\varepsilon).$
Remark F.4.
Lemma F.3 is Lemma 11 in the third arXiv version of [63], and it was further modified by Theorem 2.1 in [36] and Proposition 4 in [67]. We make some minor adjustments to our settings.
Lemma F.5.
@@
Proof of Part 1. Following Part 5 of Lemma C.1 and Eq. (1), we have:
@@
Proof of Part 2. Here, we use $R$ to denote the weight distance between arbitrary two functions $F,F^{\prime}\in{\cal F}_{\sf M,N}$, following Part 3 of Lemma C.1 and some simple algebras, $R$ is bounded by $O({\sf M}B\varepsilon)$ with $\max_{F,F^{\prime}\in{\cal F}_{\sf M,N}}\|F-F^{\prime}\|_{L^{F}({\cal X})}\leq\varepsilon$, and the space of ${\cal F}_{\sf M,N}$ is an $\sf M$-dimensional ball with radius $O({\sf M}B\varepsilon)$. Thus, for any step size $\zeta>0$, the gap is bounded by:
@@
where the first step follows from the definition of $F$ and Part 13 of Lemma C.1, the second step follows from choosing $\zeta=\varepsilon$.
@@
Theorem F.6 (Formal version of Theorem 6.2).
Let all pre-conditions hold as Corollary E.3 and choose $\varepsilon=\Theta(1/{\sf M})$. Then, with a probability at least $1-\delta$, there exists:
@@
We follow the results in Corollary E.3 ($\varepsilon\lesssim{\sf M}^{-1}$), we have:
@@
where this step follows from Part 1 of Lemma F.5 and $\xi$ is a constant.
@@
where the first step follows from Part 1 and 2 of Lemma F.5.
Result. We combine all results above and Theorem F.3, obtaining:
@@
F.2 General Scaling Law
Theorem F.7 (Formal version of Theorem 7.1).
Let all pre-conditions hold as Corollary E.3 and choose $\varepsilon=\Theta(\xi/{\sf N})$. Then, with a probability at least $1-\delta$, the following condition divides the scaling trend into two stages:
@@ F.2 General Scaling Law
where the first step follows from Eq. (F.1), the second step follows from simple algebras.
@@
Following Part 12 of Lemma C.1 and the definition of $F$, we get:
@@
Finally, we combine it with Theorem F.6 and Eq. (10):
@@
where $W(\cdot)$ is the Lambert W function.
Theorem F.8 (Single law, formal version of Theorem 7.2).
Let all pre-conditions hold as Corollary E.3. Then, with a probability at least $1-\delta$, there exists:
@@
Proof of Time-Law. This proof could be directly derived from Condition 1 in the proof of Theorem F.7.
Proof of Data-Law. This proof could be directly derived from Condition 2 in the proof of Theorem F.7.
Proof of Model-Law. Given that $\zeta\in(0,1/3]$, we choose $\varepsilon=\xi/{\sf M}^{\zeta}$, similar to Condition 2 in the proof of Theorem F.7, we choose ${\sf T}=\frac{{\sf M}^{2\zeta}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$, then we get:

### v2

@@ F.3 Approximation
Appendix G Scaling Law
G.1 Generalization Bound of Empirical Risk Minimizer
Remark G.1.
@@ G.1 Generalization Bound of Empirical Risk Minimizer
Lemma G.3 (Lemma 11 of [78], modified by Theorem 2.1 in [45] and in the setting of this paper).
@@
$\|F-F^{*}\|_{L^{F}({\cal X})}^{2}\leq 4\inf_{F^{\prime}\in{\cal F}_{\sf M,N}(\mathbb{D})}\|F-F^{*}\|_{L^{F}({\cal X})}^{2}+L{d}\cdot O(\frac{(B_{F}^{2}{+\xi^{2})}V_{({\cal F}_{\sf M,N}(\mathbb{D}),{\rm d})}(\varepsilon)}{n}+(B_{F}{+\xi)}\varepsilon).$
Remark G.4.
Lemma G.3 is Lemma 11 in the third arXiv version of [78], and it was further modified by Theorem 2.1 in [45] and Proposition 4 in [85]. We make some minor adjustments to our settings.
Lemma G.5.
@@
Proof of Part 1. Following Part 5 of Lemma D.1 and Eq. (1), we have:
@@
Proof of Part 2. Here, we use $R$ to denote the weight distance between arbitrary two functions $F,F^{\prime}\in{\cal F}_{\sf M,N}$, following Part 3 of Lemma D.1 and some simple algebras, $R$ is bounded by $O({\sf M}B\varepsilon)$ with $\max_{F,F^{\prime}\in{\cal F}_{\sf M,N}}\|F-F^{\prime}\|_{L^{F}({\cal X})}\leq\varepsilon$, and the space of ${\cal F}_{\sf M,N}$ is an $\sf M$-dimensional ball with radius $O({\sf M}B\varepsilon)$. Thus, for any step size $\zeta>0$, the gap is bounded by:
@@
where the first step follows from the definition of $F$ and Part 13 of Lemma D.1, the second step follows from choosing $\zeta=\varepsilon$.
@@
Theorem G.6 (Formal version of Theorem 5.7).
Let all pre-conditions hold as Corollary F.3 and choose $\varepsilon=\Theta(1/{\sf M})$. Then, with a probability at least $1-\delta$, there exists:
@@
We follow the results in Corollary F.3 ($\varepsilon\lesssim{\sf M}^{-1}$), we have:
@@
where this step follows from Part 1 of Lemma G.5 and $\xi$ is a constant.
@@
where the first step follows from Part 1 and 2 of Lemma G.5.
Result. We combine all results above and Theorem G.3, obtaining:
@@
G.2 General Scaling Law
Theorem G.7 (Formal version of Theorem 6.1).
Let all pre-conditions hold as Corollary F.3 and choose $\varepsilon=\Theta(\xi/{\sf N})$. Then, with a probability at least $1-\delta$, the following condition divides the scaling trend into two stages:
@@ G.2 General Scaling Law
where the first step follows from Eq. (G.1), the second step follows from simple algebras.
@@
Following Part 12 of Lemma D.1 and the definition of $F$, we get:
@@
Finally, we combine it with Theorem G.6 and Eq. (10):
@@
where $W(\cdot)$ is the Lambert W function. ∎
Theorem G.8 (Single law, formal version of Theorem 6.2).
Let all pre-conditions hold as Corollary F.3. Then, with a probability at least $1-\delta$, there exists:
@@
Proof of Time-Law. This proof could be directly derived from Condition 1 in the proof of Theorem G.7.
Proof of Data-Law. This proof could be directly derived from Condition 2 in the proof of Theorem G.7.
Proof of Model-Law. Given that $\zeta\in(0,1/3]$, we choose $\varepsilon=\xi/{\sf M}^{\zeta}$, similar to Condition 2 in the proof of Theorem G.7, we choose ${\sf T}=\frac{{\sf M}^{2\zeta}\log({\sf N}\cdot Ld/\xi^{2})}{\xi^{2}}$, then we get: